# High-dimensional analysis of ridge regression for non-identically distributed data with a variance profile

### v2

@@ 1 Introduction
$Y_{n}=X_{n}\theta_{\ast}+\varepsilon_{n},$ (1.1)
where $X_{n}$ is $n\times p$ matrix of random predictors, $\varepsilon_{n}\in{\mathbb{R}}^{n}$ is a noise vector independent of $X_{n}$ with $\mathbb{E}[\varepsilon_{n}]=0$ and $\mathbb{E}[\varepsilon_{n}\varepsilon_{n}^{t}]=\sigma^{2}I_{n}$, $\theta_{\ast}\in{\mathbb{R}}^{p}$ is a vector of unknown parameters, and $Y_{n}\in{\mathbb{R}}^{n}$ is the vector of observed responses.
Classically, the predictors are assumed to be independent and identically distributed (iid) data, meaning that the rows of the matrix $X_{n}$ are independent vectors sampled from the same probability distribution. In this paper, we propose to depart from this assumption by considering the setting where the rows of $X_{n}$ are independent but non-identically distributed. To this end, we suppose that $X_{n}$ is expressed in the following form
$X_{n}=\Upsilon_{n}\circ Z_{n},$
where $\circ$ denotes the Hadamard product between two matrices, $Z_{n}=(Z_{ij})$ has iid centered entries with variance one, and $\Upsilon_{n}=(\gamma_{ij}^{(n)})$ is a deterministic matrix. To simplify the notation, we shall sometimes write $\gamma_{ij}^{(n)}=\gamma_{ij}$ and thus drop the (possibly) dependence of $\gamma_{ij}$ on $n$. The matrix
$\Gamma_{n}=(\gamma_{ij}^{2})\in{\mathbb{R}}^{n\times p}$
governs the variance of the entries of $X_{n}$, and it is called a variance profile.
In the RMT literature, there exist various works on the analysis of the spectrum of large random matrices with a variance profile [Shl96, HLN07, ACD+21, BM20, EYY12, AEK17a]. In particular, we rely on results from [HLN07] to obtain a deterministic equivalent of the spectral distribution of a data matrix $X_{n}=\Upsilon_{n}\circ Z_{n}$ with a variance profile matrix $\Gamma_{n}$. The motivation for studying linear regression using such a variance profile is to consider the setting where one has $n$ independent pairs of observations $(Y_{i},X_{i})_{1\leq i\leq n}$ (with $X_{i}=(X_{ij})_{1\leq j\leq p}$) that are not necessarily identically distributed. Note that in the standard setting of iid data, one that
$\gamma_{ij}=\gamma_{j}\quad\mbox{for all}\quad 1\leq i\leq n,\mbox{ and }1\leq j\leq p.$
@@
We consider the high-dimensional context (with $p$ growing to infinity at a rate proportional to $n$) for which the least squares estimator is possibly not uniquely defined. Thus, we focus our analysis on the ridge regression estimator that is the minimizer of the following loss function
$\hat{\theta}_{\lambda}=\operatorname*{arg\,min}_{\theta\in{\mathbb{R}}^{p}}\frac{1}{n}\|Y_{n}-X_{n}\theta\|^{2}+\lambda\|\theta\|^{2},$
for some regularization parameter $\lambda>0$. Regardless of the ratio between $n$ and $p$, this estimator has the following explicit expression
$\hat{\theta}_{\lambda}=(X^{t}_{n}X_{n}+n\lambda I_{p})^{-1}X_{n}^{t}Y_{n}=X_{n}^{t}(X_{n}X_{n}^{t}+n\lambda I_{n})^{-1}Y_{n}.$ (1.2)
@@
To study the statistical performances of the ridge regression estimator, we analyse its predictive risk defined as
$\hat{r}_{\lambda}(X_{n})=\mathbb{E}[(\widetilde{Y}-\widetilde{X}^{t}\hat{\theta}_{\lambda})^{2}|X_{n}],$ (1.3)
where $(\widetilde{Y},\widetilde{X})\in{\mathbb{R}}\times{\mathbb{R}}^{p}$ is independent from $(Y_{n},X_{n})$ and satisfies
$\widetilde{Y}=\widetilde{X}^{t}\theta_{\ast}+\tilde{\varepsilon},\mbox{ with }\mathbb{E}[\tilde{\varepsilon}]=0,\;\mathbb{E}[\tilde{\varepsilon}^{2}]=\sigma^{2}\ \mathrm{and}\ \widetilde{X},\ \tilde{\varepsilon}\ \mathrm{independent}.$
In the above formula, $\widetilde{X}=\widetilde{S}_{p}^{1/2}\widetilde{Z}$ with $\widetilde{Z}\in{\mathbb{R}}^{p}$ a random vector with iid centered entries and variance one, and
$\widetilde{S}_{p}=\mathbb{E}[\widetilde{X}\widetilde{X}^{t}]=\mathop{\rm diag}\nolimits(\tilde{\gamma}_{1}^{2},\ldots,\tilde{\gamma}_{p}^{2})$
denotes the variance profile of $\widetilde{X}$. Note that the risk $\hat{r}_{\lambda}(X_{n})$ is conditioned on the predictors $X_{n}$, and it is thus a random variable.
Following [DW18], we focus on a random-effect hypothesis by assuming that the components of the vector $\theta_{\ast}$ are drawn independently at random. As argued in [DW18], this assumption corresponds to an average case analysis over a set of dense regression coefficients as opposed to the “sparsity hypothesis” [HTW15] or the “manifold hypothesis” [LHT23] that are other popular assumptions in high-dimensional linear regression. More precisely, the following random coefficients assumption is made throughout the paper.
@@
The vector $\theta_{\ast}$ of regression coefficients is random, independent from $X_{n}$, $\widetilde{X}$, $\varepsilon_{n}$ and $\tilde{\varepsilon}$, with $\mathbb{E}[\theta_{\ast}]=0$ and
$\mathbb{E}[\theta_{\ast}\theta_{\ast}^{t}]=\frac{\alpha^{2}}{p}I_{p}.$
The above coefficient $\alpha>0$ represents the average amount of signal strength in model (1.1). Under Assumption 1.1, the expectation in (1.3) used to define the predictive risk is thus taken with respect to both the randomness of the vector of coefficients $\theta_{\ast}$, the vector $\widetilde{X}$ and the additive noise $\varepsilon_{n}$ and $\tilde{\varepsilon}$.
@@ 1.1 Main contributions
$\hat{Y}_{\lambda}=X_{n}\hat{\theta}_{\lambda}=A_{\lambda}Y_{n},\quad\mbox{where}\quad A_{\lambda}=X_{n}(X^{t}_{n}X_{n}+n\lambda I_{p})^{-1}X_{n}^{t}.$
Then, the so-called degrees of freedom (DOF) of the estimator $\hat{\theta}_{\lambda}$ that is defined as
$\widehat{df}_{1}(\lambda)=\operatorname{Tr}[A_{\lambda}]=\operatorname{Tr}[\widehat{\Sigma}_{n}(\widehat{\Sigma}_{n}+\lambda I_{p})^{-1}],\mbox{ where }\widehat{\Sigma}_{n}=\frac{1}{n}X_{n}^{t}X_{n},$
represents the so-called effective dimension of the linear estimator $\hat{Y}_{\lambda}$. The DOF is widely used in statistics to define various criteria for model selection among a collection of estimators, see e.g. [Efr04].
Inspired by recent results from [Bac24] in the setting of iid data, a first contribution of this work is to prove the following deterministic equivalence of the DOF
@@
where
$\Sigma_{n}=\mathbb{E}[\widehat{\Sigma}_{n}]=\frac{1}{n}\mathop{\rm diag}\nolimits\left(\sum_{i=1}^{n}\gamma^{2}_{i1},\ldots,\sum_{i=1}^{n}\gamma^{2}_{ip}\right),$
and $\kappa(\lambda)$ is diagonal matrix that depends upon the regularization parameter $\lambda$ and the variance profile matrix $\Gamma_{n}$.
Throughout the paper, the meaning of the equivalence notation $A_{n}\sim B_{n}$ between two random variables is
$\lim_{n\to\infty,\;p/n\to c}|A_{n}-B_{n}|=0,\;\mbox{almost surely}.$
@@
Then, the second and main contribution of the paper is to derive a deterministic equivalent of the predictive risk $\hat{r}_{\lambda}(X_{n})$ in the case where the number of samples $n$ and the dimension $p$ tend to infinity at a proportional rate, that is $\lim\limits_{n\to+\infty}\frac{p}{n}=c>0$. This deterministic equivalent allows us to understand the influence of the ratio $c$ on the predictive risk and to also analyze the effect of the signal strength $\alpha$. We also study the convergence of the predictive risk as $\lambda$ tends to 00 to analyze the statistical properties the minimum norm least square estimator. In this setting, it appears a phenomenon arising from the curse of dimensionality that is commonly known as double descent for iid data. This phenomenon contradicts the consensus heuristic that, when a model becomes over-parameterized, then the predictive risk increases due to overfitting of the training data and the model is no longer capable of generalizing. This double descent has been thoroughly studied in the case of high-dimensional linear regression using tool from RMT, see e.g. [HMRT22, Bac24, BHX20] and references therein. In this paper, we show that it also occurs for non iid data with a variance profile. Our deterministic equivalent of the predictive risk also allows to derive the asymptotic behavior of an optimal choice for the regularization parameter and to compare it to the one obtained for iid data in [DW18].
As a third contribution, using synthetic data and various illustrative examples of variance profile, we conduct numerical experiments to verify the accuracy of our deterministic equivalent of the predictive risk using finite samples. We also investigate the similarities and differences that exist between the standard setting of iid data and the one of non-identically distributed data with a variance profile.
For example, if the variance profile is assumed to be quasi doubly stochastic in the sense that
$\frac{1}{n}\sum_{j=1}^{p}\gamma_{ij}^{2}=\frac{p}{n},\mbox{ for all }1\leq i\leq n,\quad\mbox{and}\quad\frac{1}{n}\sum_{i=1}^{n}\gamma_{ij}^{2}=1,\mbox{ for all }1\leq j\leq p,$
@@
$\lim_{n\to\infty,\;p/n\to c}\hat{r}_{0}(X_{n})=\left\{\begin{array}[]{ccc}\sigma^{2}\frac{c}{1-c}+\sigma^{2}&\mbox{if}&c<1\\ \alpha^{2}\left(1-\frac{1}{c}\right)+\sigma^{2}\frac{1}{c-1}&\mbox{if}&c>1\end{array}\right.,\;\mbox{almost surely}.$ (1.5)
@@
$\Gamma_{n}=\left[\begin{array}[]{cc}\gamma_{1}^{2}\mathbbm{1}_{n/4}\mathbbm{1}_{p/4}^{t}&\gamma_{2}^{2}\mathbbm{1}_{n/4}\mathbbm{1}_{3p/4}^{t}\\ \gamma_{2}^{2}\mathbbm{1}_{3n/4}\mathbbm{1}_{p/4}^{t}&\gamma_{1}^{2}\mathbbm{1}_{3n/4}\mathbbm{1}_{3p/4}^{t}\end{array}\right],$

### v3

@@ 1 Introduction
$y=x^{\top}\beta_{\ast}+\varepsilon,$ (1.1)
where $x\in\mathbb{R}^{p}$ is a vector of random predictors, $\varepsilon\in{\mathbb{R}}$ is a noise vector independent of $x$ with $\mathbb{E}[\varepsilon]=0$ and $\mathbb{E}[\varepsilon^{2}]=\sigma^{2}>0$, $\beta_{\ast}\in{\mathbb{R}}^{p}$ is a vector of unknown parameters, and $y\in{\mathbb{R}}$ is the observed response. Let consider a training sample $(x_{1},y_{1}),...,(x_{n},y_{n})$ following this linear model, that is for $i=1,...,n$, one has that $y_{i}=x_{i}^{\top}\beta_{\ast}+\varepsilon_{i}$ where the noise term $\varepsilon_{i}$ is defined as above. Then one obtain from (1.1) that $Y_{n}=X_{n}\beta_{\ast}+\varepsilon_{n},$ where $X_{n}=(x_{1}|\dots|x_{n})^{\top}\in\mathbb{R}^{n\times p}$, $Y_{n}=(y_{1}|\dots|y_{n})^{\top}\in\mathbb{R}^{n}$ and $\varepsilon_{n}=(\varepsilon_{1}|\dots|\varepsilon_{n})^{\top}\in\mathbb{R}^{n}$.
Classically, the predictors are assumed to be independent and identically distributed (iid) data, meaning that the rows of the matrix $X_{n}$ are independent vectors sampled from the same probability distribution. In this paper, we propose to depart from this assumption by considering the setting where the rows of $X_{n}$ are independent but non-identically distributed. To this end, we suppose that $X_{n}$ is expressed in the following form $X_{n}=\Upsilon_{n}\circ Z_{n},$ where $\circ$ denotes the Hadamard product between two matrices, $Z_{n}=(Z_{ij})$ has iid centered entries with variance one, and $\Upsilon_{n}=(\gamma_{ij}^{(n)})$ is a deterministic matrix. To simplify the notation, we shall sometimes write $\gamma_{ij}^{(n)}=\gamma_{ij}$ and thus drop the (possibly) dependence of $\gamma_{ij}$ on $n$. The matrix $\Gamma_{n}=(\gamma_{ij}^{2})\in{\mathbb{R}}^{n\times p}$ governs the variance of the entries of $X_{n}$, and it is called a variance profile.
In the RMT literature, there exist various works on the analysis of the spectrum of large random matrices with a variance profile [Shl96, HLN07, ACD+21, BM20, EYY12, AEK17a]. In particular, we rely on results from [HLN07] to obtain a deterministic equivalent of the spectral distribution of a data matrix $X_{n}=\Upsilon_{n}\circ Z_{n}$ with a variance profile matrix $\Gamma_{n}$. The motivation for studying linear regression using such a variance profile is to consider the setting where one has $n$ independent pairs of observations $(Y_{i},X_{i})_{1\leq i\leq n}$ (with $X_{i}=(X_{ij})_{1\leq j\leq p}$) that are not necessarily identically distributed. Note that in the standard setting of iid data, one has that $\gamma_{ij}=\gamma_{j}\quad\mbox{for all}\quad 1\leq i\leq n,\mbox{ and }1\leq j\leq p.$
@@
As an application, our methodology offers a novel tool for analyzing data that arises from a mixture model that is a classical framework in machine learning, when the data are sampled from multiple underlying subpopulations or classes. Indeed, consider latent class variables $C_{1},\ldots,C_{K}$, which determine the class membership of each feature vector $x_{i}$. Formally, for each $i$ and each class $k\in\{1,\ldots,K\}$, we assume that the latent class variable $C_{i}$ follows a categorical distribution: $\forall 1\leq i\leq n,\quad\mathbb{P}(C_{i}=k)=\pi_{k},$ Within each class, the random predictor $x_{i}$ is then assumed to follow a specific covariance structure. Conditional on $\{C_{i}=k\}$, we model the predictors as $x_{i}=S_{k}^{1/2}x_{i}^{\prime},$ where $S_{k}=\text{diag}(s_{k,1}^{2},\ldots,s_{k,p}^{2})$ is a diagonal matrix that characterizes the covariance structure of the predictors within class $k$. The variances $s_{k,j}^{2}$ of the predictors vary across classes, reflecting potential heterogeneity in the data. Then, given the class labels $C_{1},\ldots,C_{n}$, the resulting matrix of predictors $X_{n}$ exhibits a variance profile governed by the matrix $\Gamma_{n}=(s_{C_{i},j}^{2})\in\mathbb{R}^{n\times p},$ where each entry $s_{C_{i},j}^{2}$ corresponds to the variance of the $j$-th feature for the $i$-th feature vector, determined by its class membership $C_{i}$. Hence, our approach allows to investigate the double descent phenomenon in the context of high-dimensional ridge regression when the data follow a mixture model, an aspect that has not been previously explored in the literature, to the best of our knowledge.
We consider the high-dimensional context (with $p$ growing to infinity at a rate proportional to $n$) for which the least squares estimator is possibly not uniquely defined. Thus, we focus our analysis on the ridge regression estimator that is the minimizer of the following loss function $\hat{\theta}_{\lambda}=\operatorname*{arg\,min}_{\theta\in{\mathbb{R}}^{p}}\frac{1}{n}\|Y_{n}-X_{n}\theta\|^{2}+\lambda\|\theta\|^{2},$ for some regularization parameter $\lambda>0$. Regardless of the ratio between $n$ and $p$, this estimator has the following explicit expression
$\hat{\theta}_{\lambda}=(X^{\top}_{n}X_{n}+n\lambda I_{p})^{-1}X_{n}^{\top}Y_{n}=X_{n}^{\top}(X_{n}X_{n}^{\top}+n\lambda I_{n})^{-1}Y_{n}.$ (1.2)
@@
To study the statistical performances of the ridge regression estimator, we analyse its predictive risk, denoted $\hat{r}_{\lambda}^{test}(X_{n})$, and its train risk, denoted $\hat{r}_{\lambda}^{train}(X_{n})$, defined as
$\hat{r}_{\lambda}^{test}(X_{n})=\mathbb{E}[(\tilde{y}-\tilde{x}^{\top}\hat{\theta}_{\lambda})^{2}|X_{n}],$ (1.3)
$\hat{r}_{\lambda}^{train}(X_{n})=\frac{1}{n}\mathbb{E}[\parallel Y_{n}-X_{n}\hat{\theta}_{\lambda}\parallel^{2}|X_{n}],$ (1.4)
where $(\tilde{y},\tilde{x})\in{\mathbb{R}}\times{\mathbb{R}}^{p}$ is independent from $(Y_{n},X_{n})$ and satisfies
$\tilde{y}=\tilde{x}^{\top}\beta_{\ast}+\tilde{\varepsilon},\mbox{ with }\mathbb{E}[\tilde{\varepsilon}]=0,\;\mathbb{E}[\tilde{\varepsilon}^{2}]=\sigma^{2}\ \mathrm{and}\ \widetilde{X},\ \tilde{\varepsilon}\ \mathrm{independent}.$
In the above formula, $\tilde{x}=\widetilde{S}_{p}^{1/2}\widetilde{z}$ with $\widetilde{z}\in{\mathbb{R}}^{p}$ a random vector with iid centered entries and variance one, and $\widetilde{S}_{p}=\mathbb{E}[\tilde{x}\tilde{x}^{\top}]=\mathop{\rm diag}\nolimits(\tilde{\gamma}_{1}^{2},\ldots,\tilde{\gamma}_{p}^{2})$ denotes the variance profile of $\tilde{x}$. Note that the risk $\hat{r}_{\lambda}^{test}(X_{n})$ is conditioned on the predictors $X_{n}$, and it is thus a random variable.
Following [DW18], we focus on a random-effect hypothesis by assuming that the components of the vector $\beta_{\ast}$ are drawn independently at random. As argued in [DW18], this assumption corresponds to an average case analysis over a set of dense regression coefficients as opposed to the “sparsity hypothesis” [HTW15] or the “manifold hypothesis” [LHT23] that are other popular assumptions in high-dimensional linear regression. The following assumptions are made throughout the paper, and they are used to derive deterministic equivalents of the training and predictive risks.
@@
The vector $\beta_{\ast}$ of regression coefficients is random, independent from $X_{n}$, $\widetilde{x}$, $\varepsilon_{n}$ and $\tilde{\varepsilon}$, with $\mathbb{E}[\beta_{\ast}]=0$ and $\mathbb{E}[\beta_{\ast}\beta_{\ast}^{\top}]=\frac{\alpha^{2}}{p}I_{p}.$
Assumption 1.2.
$\exists\ \delta>0$ s.t. $\mathbb{E}[|Z_{ij}|^{4+\delta}],\mathbb{E}[|\widetilde{z}_{j}|^{4+\delta}]<+\infty,$ $\forall\ 1\leq i\leq n$ and $1\leq j\leq p$.
Assumption 1.3.
$\exists\ \gamma_{\mathrm{max}}>0$ s.t. $\underset{n\geq 1}{\mathrm{sup}}\underset{\begin{subarray}{c}1\leq i\leq n\\ 1\leq j\leq p\end{subarray}}{\mathrm{max}}\{|\gamma_{ij}^{(n)}|,|\tilde{\gamma}_{j}^{(n)}|\}<\gamma_{\mathrm{max}}.$
Assumption 1.4.
$\exists\ \gamma_{\mathrm{min}}>0$ s.t. $\forall n\geq 1,\underset{\begin{subarray}{c}1\leq i\leq n\\ 1\leq j\leq n\end{subarray}}{\mathrm{min}}\{|\gamma_{ij}^{(n)}|,|\tilde{\gamma}_{j}^{(n)}|\}\geq\gamma_{\mathrm{min}}.$
Assumption 1.5.
$\exists\ d_{\ast}>0$ s.t. $|\frac{p}{n}-1|\geq d_{\ast}$ for any values of $n$ and $p$.
The coefficient $\alpha>0$ represents the average amount of signal strength in model (1.1). Under Assumption 1.1, the expectation in (1.3) used to define the predictive risk is thus taken with respect to both the randomness of the vector of coefficients $\beta_{\ast}$, the vector $\tilde{x}$ and the additive noise $\varepsilon_{n}$ and $\tilde{\varepsilon}$. Assumptions 1.2 to 1.5 ensures that the spectrum of $\frac{1}{n}X_{n}^{\top}X_{n}$ well behave in the high dimensional framework, that is whenever $\min(n,p)\to+\infty$ and $\frac{p}{n}\to c>0$. Indeed, it is proved in [HLN07] that Assumptions 1.2 and 1.3 guarantees that the empirical singular value distribution of $X_{n}$ converges, in the high dimensional regime, to a distribution solution of a fixed point equation. Note that these assumptions are not limited to the case of random Gaussian data. On the other hand, [AEK17b] proves that Assumptions 1.4 and 1.5 ensures that the spectrum of $\frac{1}{n}X_{n}^{\top}X_{n}$ is bounded away from 00. Note that we only use this condition to study the behavior of the risks in the ridge (less) framework, that is whenever $\lambda=0$. Hence Assumptions 1.4 and 1.5 can be ignored as long as $\lambda\neq 0$.
@@ 1.1 Main contributions
$\hat{Y}_{\lambda}=X_{n}\hat{\theta}_{\lambda}=A_{\lambda}Y_{n},\quad\mbox{where}\quad A_{\lambda}=X_{n}(X^{\top}_{n}X_{n}+n\lambda I_{p})^{-1}X_{n}^{\top}.$
Then, the degrees of freedom (DOF) of the estimator $\hat{\theta}_{\lambda}$, that is defined as
$\widehat{df}_{1}(\lambda)=\operatorname{Tr}[A_{\lambda}]=\operatorname{Tr}[\widehat{\Sigma}_{n}(\widehat{\Sigma}_{n}+\lambda I_{p})^{-1}],\mbox{ where }\widehat{\Sigma}_{n}=\frac{1}{n}X_{n}^{\top}X_{n},$
represents the so-called effective dimension of the linear estimator $\hat{Y}_{\lambda}$. The DOF is widely used in statistics to define various criteria for model selection among a collection of estimators, see e.g. [Efr04]. Inspired by recent results from [Bac24] in the setting of iid data, a first contribution of this work is to prove the following deterministic equivalence of the DOF
@@
where $\Sigma_{n}=\mathbb{E}[\widehat{\Sigma}_{n}]=\frac{1}{n}\mathop{\rm diag}\nolimits\left(\sum_{i=1}^{n}\gamma^{2}_{i1},\ldots,\sum_{i=1}^{n}\gamma^{2}_{ip}\right),$ and $\kappa(\lambda)$ is diagonal matrix that depends upon the regularization parameter $\lambda$ and the variance profile matrix $\Gamma_{n}$.
Throughout the paper, the meaning of the equivalence notation $A_{n}\sim B_{n}$ between two random variables is $\lim_{n\to\infty,\;p/n\to c}|A_{n}-B_{n}|=0,\;\mbox{almost surely}.$
@@
Then, the second and main contribution of the paper is to derive a deterministic equivalent of the predictive risk $\hat{r}_{\lambda}^{test}(X_{n})$ and the training risk $\hat{r}_{\lambda}^{train}(X_{n})$ in the case where the number of samples $n$ and the dimension $p$ tend to infinity at a proportional rate, that is $\lim\limits_{n\to+\infty}\frac{p}{n}=c>0$. This deterministic equivalent allows us to understand the influence of the ratio $c$ on the predictive risk and to also analyze the effect of the signal strength $\alpha$. We also study the convergence of the predictive risk as $\lambda$ tends to 00 to analyze the statistical properties the minimum norm least square estimator. In this setting, it appears a phenomenon arising from the curse of dimensionality that is commonly known as double descent for iid data. This phenomenon contradicts the consensus heuristic that, when a model becomes over-parameterized, then the predictive risk increases due to overfitting of the training data and the model is no longer capable of generalizing. This double descent has been thoroughly studied in the case of high-dimensional linear regression using tool from RMT, see e.g. [HMRT22, Bac24, BHX20] and references therein. In this paper, we show that it also occurs for non iid data with a variance profile. Our deterministic equivalent of the predictive risk also allows to derive the asymptotic behavior of an optimal choice for the regularization parameter and to compare it to the one obtained for iid data in [DW18].
As a third contribution, using synthetic data and various illustrative examples of variance profile, we conduct numerical experiments to verify the accuracy of our deterministic equivalent of the predictive risk using finite samples.
We also investigate the similarities and differences that exist between the standard setting of iid data and the one of non-identically distributed data with a variance profile. For example, if the variance profile is assumed to be quasi doubly stochastic in the sense that
$\frac{1}{n}\sum_{j=1}^{p}\gamma_{ij}^{2}=\frac{p}{n},\mbox{ for all }1\leq i\leq n,\quad\mbox{and}\quad\frac{1}{n}\sum_{i=1}^{n}\gamma_{ij}^{2}=1,\mbox{ for all }1\leq j\leq p,$ (1.6)
@@
$\lim_{n\to\infty,\;p/n\to c}\hat{r}_{0}^{test}(X_{n})=\left\{\begin{array}[]{ccc}\sigma^{2}\frac{c}{1-c}+\sigma^{2}&\mbox{if}&c<1\\ \alpha^{2}\left(1-\frac{1}{c}\right)+\sigma^{2}\frac{1}{c-1}&\mbox{if}&c>1\end{array}\right.,\;\mbox{almost surely}.$ (1.7)
@@
$\Gamma_{n}=\left[\begin{array}[]{cc}\gamma_{1}^{2}\mathbbm{1}_{n/4}\mathbbm{1}_{p/4}^{\top}&\gamma_{2}^{2}\mathbbm{1}_{n/4}\mathbbm{1}_{3p/4}^{\top}\\ \gamma_{2}^{2}\mathbbm{1}_{3n/4}\mathbbm{1}_{p/4}^{\top}&\gamma_{1}^{2}\mathbbm{1}_{3n/4}\mathbbm{1}_{3p/4}^{\top}\end{array}\right],$